OpenStudy (anonymous):

Can someone please explain how I would do this? Arrange the functions x^1/2, x^2, x^3, and 2^x by relative sizes of their y-values a.On (0,1)

OpenStudy (loser66):

just plug x =0 into each of the functions, you can calculate the value of them, then arrange them in increasing or decreasing order for example, for \[y= x^{\frac{1}{2}}\] plug x =0, you have \[0^\frac{1}{2}=0 \]

OpenStudy (loser66):

and then to other functions. for x =0 you have 4 values of functions, arrange them. for x =1 do the same/ hihihi.... I don't know whether my understanding is right or wrong, just give out what I think.

OpenStudy (anonymous):

I'll try, thank you!

OpenStudy (loser66):

let hear other 's opinion. @e.mccormick

OpenStudy (e.mccormick):

https://www.desmos.com/calculator/vjm827rwpo Depends on the evaluation point, but yes, they want you to set the x, get the y, and compare those ys.

OpenStudy (loser66):

bingo, thank you e.mccormick!!!

OpenStudy (anonymous):

So 1=0^1/2, 1=0^2, 1=0^3, 1=2^0?

OpenStudy (e.mccormick):

(0,1) is an interval. because it says on.

OpenStudy (e.mccormick):

"On (0,1)" means \(0<x<1\)

OpenStudy (e.mccormick):

If you look at the point (1,1) on the graph I linked, you can see why they said on that interval.

OpenStudy (campbell_st):

a really easy way to do this question is to graph all 4 curves  and the link below is to the best free graphing & geometry package available http://www.geogebra.org/cms/en/